# Authorship Verification based on the Likelihood Ratio of Grammar Models

### v1

@@ 5.1.12 $\mathcal{C}_{\mathrm{AllNews}}$
The $\mathcal{C}_{\mathrm{AllNews}}$ corpus consists of 4,438 news articles written by 2,219 authors. The articles stem from the large well-known All the News 2.0 dataset [^12] The dataset can be accessed at https://components.one/datasets/all-the-news-2-news-articles-dataset, which consists of 2.7 million news articles and essays from 27 American news agencies and podcasting companies. From these, we limit ourselves to a single source, namely Reuters, which contains the largest number of news articles (840,094 in total). However, since not all articles were suitable for the creation of $\mathcal{C}_{\mathrm{AllNews}}$, it was necessary to preprocess the initial Reuters subset accordingly. First, we discarded all authors who wrote only a single article, as otherwise there is no suitable possibility to construct two verification cases for each author. Next, we sorted all articles of each author by their publication year and kept only four articles (the two oldest and the two newest). The main reasons for this were firstly to maximize the stylistic variation of each author and secondly to ensure sufficient length of the resulting documents, which in some cases were massively shortened by the pre-processing. Based on the four documents per author, the pre-processing was performed as follows. First, we removed all quotations (which are very common due to the nature of news articles) to mitigate the degree of stylistic distortion. Next, we removed headlines, sentences that consisted of more than 40 percent named entities and/or numeric tokens (e. g., dates, amounts of money, percentages, etc.), and article signatures that contained names of people who reported, wrote, and edited the articles. Finally, we normalized apostrophes (e. g., $\rightarrow$”) and replaced multiple consecutive spaces with a single space. The pre-processing procedure severely impaired the coherence of the resulting documents, as the natural order of many sentences was no longer given and some sentences also resulted in a truncated form. The impaired coherence can be seen as a challenge of this corpus, as AV methods that use implicit features (e. g., character/word $n$-grams) may capture “artificial features” that occur across sentence boundaries. Another challenge of $\mathcal{C}_{\mathrm{AllNews}}$ is that the documents often have been edited by other authors (e. g., the editor-in-chief) before online publication, so the original writing style of the articles may be distorted to some extent. Unlike the Stack Exchange network, All the News 2.0 does not provide metadata indicating which specific parts of the original article have been edited. Therefore, it was not possible for us to remove these afterwards.
@@ 5.2 Topic-Masking via POSNoise
Authorship verification constitutes a similarity detection problem in which the subject of the similarity determination is the language of the author rather than other document aspects such as the topic [42]. However, a large proportion of existing AV methods including [1, 17, 19, 22, 69, 70, 72, 80, 88] use features that are directly influenced by the topic and thus run the risk of biasing their predictions accordingly. To counteract the problem of biased predictions regarding LambdaG, which also employs topic-affected features, we apply topic masking to all documents in our corpora. Topic masking is an approach to pre-process texts in such a way that topic-dependent text units (e. g., words, phrases, etc.) are masked by appropriate placeholders, while grammatical text units (e. g., determiners, conjunctions, contractions, prepositions, pronouns, auxiliary verbs, quantifiers, transitional phrases, etc.) are retained within the texts. Of the few topic masking approaches published in the literature, we chose POSNoise presented by Halvani and Graner [42], which proved to outperform the previously published topic masking approach called TextDistortion [103] in 34 out of 42 cases, with an increase in accuracy of up to 10%.
@@
Table 5: Comparison between the original texts and their resulting topic-masked representations generated by POSNoise [42, Algorithm: 1]. The semantics of the placeholders can be looked up in [42, Table: 2].
5.3 N-Gram Language Modelling with Kneser-Ney Smoothing
@@ 5.3.2 Prefix and Suffix Counts
where ||$\cdot$|| ⋅ | specifies the cardinality of a set. In plain words, these functions provide the number of different prefixes or suffixes a $n$-gram has been observed with exactly, or at least, $r$ times. Notice that all definitions automatically extend to the case where $g=\emptyset$, and we will use the shorthand notations $N_{x}(\bullet)\equiv N_{x}(\emptyset\bullet)$ and $N_{x}(\bullet\bullet)\equiv N_{x}(\bullet\emptyset\bullet)$. The juxtaposition of the empty 00-gram to other $n$-grams is understood as no operation, e. g., $t\mbox{}\emptyset t\mbox{}^{\prime}\equiv t\mbox{}t\mbox{}^{\prime}$.
@@ 5.3.3 N-gram Language Models
where, in turn, the probability of a sentence $S\mbox{}_{i}=t\mbox{}_{i,1}t\mbox{}_{i,2}\cdots t\mbox{}_{i,n_{i}}$ is:
$P(S\mbox{}_{i})=\prod_{j=1}^{n_{i}+1}P(t\mbox{}_{i,j}|t\mbox{}_{i,j-N+1}t\mbox{}_{i,j-N+2}\cdots t\mbox{}_{i,j-1})$ (13)
and we make the conventions $t\mbox{}_{i,n_{i}+1}=\langle\text{EOS}\rangle$, $t\mbox{}_{i,-1}=t\mbox{}_{i,-2}=\cdots=t\mbox{}_{i,-N+1}=\langle\text{BOS}\rangle$. A detail that is often overlooked in the literature is that the probability in Eq. (12) is conditional on the number $m$ of sentences that compose the document $\mathcal{D}_{m}$, as reflected by our notation.
@@ 5.3.4 Kneser-Ney Smoothing
The functions $\alpha$ and $\gamma$ are defined by:
@@
Finally, $D(1),\ D(2),\,\dots$ are free parameters satisfying $10<D(k)<$ 1, whose optimal values can be estimated from data. The parameter space is usually made finite by assuming $D(k)=D(r)$ for some $r$ and $k\geq r$, in which case we can rewrite the infinite sum in (17) as:
@@ 5.4 The Likelihood Ratio Framework
  • the Defence Hypothesis ($H_{d}$): The disputed document ($\mathcal{D}_{\mathcal{U}}$) was not written by the author of the known document(s) (the defendant, $\mathcal{A}$), but was written by someone else, $\neg\mathcal{A}$
@@
It is important to note that the forensic expert cannot estimate and legally must not refer to the posterior odds. They cannot estimate it logically because they are not privy of the trier-of-fact’s belief regarding the hypotheses and do not have access to other pieces of evidence presented prior to the linguistic text evidence. Legally, they must not refer to it because the posterior odds represents the ultimate issue of being guilty or innocence [73, p. 96] and this determination is a specific responsibility reserved only for the trier-of-fact. The odds form of Bayes’ Theorem explicitly outlines the role of the forensic scientist. In the Likelihood Ratio framework, their responsibility lies in estimating the strength of evidence in the form of a likelihood ratio. As indicated by the rightmost term of Equation 20, the likelihood ratio is the ratio between the probability of the evidence given $H_{p}$ and the probability of the same evidence given $H_{d}$. The Likelihood Ratio serves as a gradient, quantifying how much more likely the evidence is given one hypothesis over the other [35]. When the LR is greater than one, it provides stronger support for $H_{p}$, whereas when it is smaller than one, it offers stronger support for $H_{d}$. The LR framework is now accepted as a logically and legally correct framework for interpreting forensic evidential analyses [3, 94], having been studied for various evidential types [15, 16, 29, 77, 81, 113]. The LR framework has also been supported by the relevant scientific and professional associations [2, 85, 23, 92, 6, 84].
@@ 5.4.1 The Log-Likelihood Ratio Cost ($C_{\text{llr}}\mbox{}$)
In numerous source-identification tasks, the system’s performance is commonly evaluated in terms of identification accuracy and/or identification error rate. Metrics such as precision, recall, and equal error rate (EER) are typical in this regard. While these metrics offer a quick overview of performance and are intuitively easy to comprehend, they are not suitable for assessing likelihood ratio-based inference systems [78, p. 93].
@@
Primarily based on categorical decisions with specific thresholds, these metrics are inadequate for evaluating such systems. Firstly, metrics relying on categorical thresholding implicitly assess how effectively the system correctly identifies the source of evidence. In other words, these metrics implicitly pertain to the strength of the hypothesis rather than the strength of the evidence [78, p. 93]. Secondly, categorical metrics fail to capture the gradient nature of likelihood ratios [78, p. 93]. For instance, in a same-author comparison, if one system returns $\log_{10}(\text{LR})=3$ while another yields $\log_{10}(\text{LR})=2$, the former is superior as it supports the correct hypothesis more strongly. However, using a categorical metric would assess both systems in the same way, disregarding the differing strengths of evidence supporting the correct hypothesis.
@@
The conventional metric employed for evaluating the performance of a system within the Likelihood Ratio Framework is the log-likelihood ratio cost ($C_{\text{llr}}\mbox{}$). The computation of the $C_{\text{llr}}\mbox{}$ involves utilising a set of LRs for same-source scenarios (e.g., same-author, Y) and a set for different-source scenarios (e. g., different-author, N). The computation is performed using the following equation [20, 91, 111]
@@
$C_{\text{llr}}\mbox{}=\frac{1}{2}\left(\left[\frac{1}{n}\sum_{i}^{n}\log_{2}\left(1+\frac{1}{LR_{i}}\right)\right]+\left[\frac{1}{m}\sum_{j}^{m}\log_{2}\bigg{(}1+LR_{j}\bigg{)}\right]\right)$ (21)
where, $n$ and $m$ are the number of Y- and N-cases, and $LR_{i}$ and $LR_{j}$ are the LRs estimated for the Y- and N-cases, respectively. $C_{\text{llr}}\mbox{}$ considers the magnitude of calculated LR values, imposing appropriate penalties based on their values. Within the $C_{\text{llr}}\mbox{}$ metric, LRs supporting counter-factual hypotheses, or contrary-to-fact LRs (with $\log_{10}(\text{LR})<0$ for Y-problems and $\log_{10}(\text{LR})>0$ for N-problems), incur significant penalties. The extent of the penalty is proportional to the deviation of the LRs from unity. Optimal performance is attained when $C_{\text{llr}}\mbox{}$ equals 0, and performance gradually diminishes as $C_{\text{llr}}\mbox{}$ approaches and surpasses 1. Therefore, a lower $C_{\text{llr}}\mbox{}$ value corresponds to better performance.
@@
$C_{\text{llr}}\mbox{}$ serves as an assessment metric for the overall performance of an LR-based inference system and can be decomposed into two distinct metrics: $C_{\text{llr}}^{\text{min}}\mbox{}$ and $C_{\text{llr}}^{\text{cal}}\mbox{}$. The former evaluates the discrimination performance and the latter the calibration performance of the system. The $C_{\text{llr}}^{\text{min}}\mbox{}$ is determined by computing the losses for the optimised LRs through the non-parametric pool-adjacent-violators algorithm. The disparity between $C_{\text{llr}}\mbox{}$ and $C_{\text{llr}}^{\text{min}}\mbox{}$ is expressed as $C_{\text{llr}}^{\text{cal}}\mbox{}$; in other words, $C_{\text{llr}}\mbox{}$ equals the sum of discrimination loss ($C_{\text{llr}}^{\text{min}}\mbox{}$) and calibration loss ($C_{\text{llr}}^{\text{cal}}\mbox{}$). More in-depth descriptions of $C_{\text{llr}}\mbox{}$, encompassing both $C_{\text{llr}}^{\text{min}}\mbox{}$ and $C_{\text{llr}}^{\text{cal}}\mbox{}$, can be found in work by Brümmer and du Preez [20], Drygajlo et al. [34], and Meuwly et al. [75]. The $C_{\text{llr}}\mbox{}$ has already been effectively utilized for evaluating LR-based forensic text comparison systems [49, 50, 51, 82].
5.5 SiamBERT: Neural Net Baseline
Recognising the paradigm shift towards neural network techniques that is also evident in the research field of AV, we consider it important to compare LambdaG with contemporary methods utilizing deep learning techniques. For this reason, we incorporate in our evaluation a baseline method based on a Siamese BERT neural network and functionally very similar to the Sentence-BERT architecture [93] and Ishihara’s et al [53] method, in which likelihood-ratios were estimated via cosine similarity and probabilistic linear discrimination analysis. We call this method from here on SiamBERT. Note that, in our experiments, we use the BERT-Large, Cased model with its corresponding tokenizer [30], which are pretrained on English text and considers character casing. Given a document $\mathcal{D}$, we tokenize it into a sequence of tokens. Since the BERT model we are using is only able to handle up to 512 tokens, we split the sequence of tokens into a set of adjacent chunks $\varsigma_{1},\ldots,\varsigma_{m}$, where $m$ is the number of chunks. Each $\varsigma_{i}$ comprises 512 tokens, except for $\varsigma_{m}$, which may comprise less. These chunks are then embedded as feature vectors $F_{1},\ldots,F_{m}$, by encoding each $\varsigma_{i}$ with the BERT model and taking the average of the resulting embedding vectors over the tokens in a chunk. Finally, the feature vectors are then again averaged, resulting in a final feature vector $F$, which represents the document $\mathcal{D}$ in an embedding space. Given a single verification case, we consider an unknown document $\mathcal{D}_{\mathcal{U}}$ and a known document $\mathcal{D}_{\mathcal{A}}$ for SiamBERT. Both are then embedded as described above, resulting in two feature vectors $F_{\mathcal{U}}$ and $F_{\mathcal{A}}$. Based on these vectors, we then calculate their cosine similarity $s$. If $s>\theta$, where $\theta$ is a threshold, we assume $\mathcal{U}=\mathcal{A}$ (i. e., same authorship). To train SiamBERT, we further split the training set $\mathcal{C_{\,\textrm{train}}}$ into $\mathcal{C_{\,\textrm{train}*}}$ and $\mathcal{C_{\,\textrm{val}}}$ based on a 80/20 ratio, in order to optimize hyper-parameters. We trained the parameters of the BERT model on $\mathcal{C_{\,\textrm{train}}}$ with MSE loss function (calculating the mean squared error between the cosine similarity and the binary label, where 1 represents a Y case), the AdamW optimizer, a batch size of $4$ and a learning rate of $0.00005$ for $100$ epochs (scheduled with a linear warmup in the first $10$ epochs and a cosine decay to zero in the remaining). In order to optimize the hyper-parameter epoch count, we measured the AUC on each epoch of the corresponding model on $\mathcal{C_{\,\textrm{val}}}$ and selected the model with the best AUC. Finally, we calculated the cosine similarities of the verification cases of $\mathcal{C_{\,\textrm{train}}}$ using the selected model and determined the optimal threshold $\theta$, which would yield the best accuracy on $\mathcal{C_{\,\textrm{train}}}$.
6 Appendix A: Qualitative Data Exploration

### v2

@@ 6.1.12 $\mathcal{C}_{\mathrm{AllNews}}$
The $\mathcal{C}_{\mathrm{AllNews}}$ corpus consists of 4,438 news articles written by 2,219 authors. The articles stem from the large well-known portal All the News 2.0 [35], which consists of 2.7 million news articles and essays from 27 American news agencies and podcasting companies. From these, we limit ourselves to a single source, namely Reuters, which contains the largest number of news articles (840,094 in total). However, since not all articles were suitable for the creation of $\mathcal{C}_{\mathrm{AllNews}}$, it was necessary to preprocess the initial Reuters subset accordingly. First, we discarded all authors who wrote only a single article, as otherwise there is no suitable possibility to construct two verification cases for each author. Next, we sorted all articles of each author by their publication year and kept only four articles (the two oldest and the two newest). The main reasons for this were, firstly, to maximize the stylistic diversity of each author and, secondly, to ensure a sufficient length of the resulting documents, which in some cases were heavily shortened due to the pre-processing. Based on the four documents per author, the pre-processing was carried out as follows. First, we removed all quotations (which are very common due to the nature of news articles) to mitigate the degree of stylistic distortion. Next, we removed headlines, sentences that consisted of more than 40 percent named entities and/or numeric tokens (e. g., dates, amounts of money, percentages, etc.), and article signatures that contained names of people who reported, wrote, and edited the articles. Finally, we normalized apostrophes (e. g., $\rightarrow$”) and replaced multiple consecutive spaces with a single space. The pre-processing procedure severely impaired the coherence of the resulting documents, as the natural order of many sentences was no longer given and some sentences also resulted in a truncated form. The impaired coherence can be seen as a challenge of this corpus, as AV methods that use implicit features (e. g., character/word $n$-grams) may capture “artificial features” that occur across sentence boundaries. Another challenge of $\mathcal{C}_{\mathrm{AllNews}}$ is that the documents often have been edited by other authors (e. g., the editor-in-chief) before online publication, so the original writing style of the articles may be distorted to some extent. Unlike the Stack Exchange network, All the News 2.0 does not provide metadata indicating which specific parts of the original article have been edited. Therefore, it was not possible for us to remove these retrospectively.
@@ 6.2 Topic-Masking via POSNoise
Figure 5: The POSNoise algorithm. Details on the notation can be found in [53].
Authorship Verification constitutes a similarity detection problem in which the subject of the similarity determination is the language of the author rather than other document aspects such as the topic [53]. However, a large number of existing AV methods including [1, 23, 25, 29, 83, 84, 86, 95, 102] use features that are directly influenced by the topic and thus run the risk of biasing their predictions accordingly. To counteract the problem of biased predictions regarding LambdaG, which also employs topic-affected features, we apply topic masking to all involved documents (see Algorithm 1). Topic masking is an approach to pre-process texts in such a way that topic-dependent text units such as words or phrases are masked by appropriate placeholders, while grammatical text units (e. g., determiners, conjunctions, contractions, prepositions, pronouns, auxiliary verbs, quantifiers, transitional phrases, etc.) are retained within the texts. Of the few topic masking approaches published in the literature, we chose POSNoise presented by Halvani and Graner [53], which proved to outperform the previously published topic masking approach called TextDistortion [119] in 34 out of 42 cases, with an increase in Accuracy of up to 10%.
Category Examples
[…]
Table 6: All categories of topic agnostic words and phrases whitelisted by POSNoise with some examples.
@@
Table 7: Comparison between the original texts and their resulting topic-masked representations generated by POSNoise [53, Algorithm: 1]. The semantics of the placeholders ©, Ø, @, # and § can be looked up in [53, Table: 2].
6.3 N-Gram Language Modeling with Kneser-Ney Smoothing
@@ 6.3.2 Prefix and Suffix Counts
where ||$\cdot$|| ⋅ | specifies the cardinality of a set. In plain words, these functions provide the number of different prefixes or suffixes a $n$-gram has been observed with exactly, or at least, $r$ times. Notice that all definitions automatically extend to the case where $g=\emptyset$, and we will use the shorthand notations $N_{r}(\bullet)\equiv N_{r}(\emptyset\bullet)$ and $N_{r}(\bullet\bullet)\equiv N_{r}(\bullet\emptyset\bullet)$. The juxtaposition of the empty 00-gram to other $n$-grams is understood as no operation, e. g., $t\mbox{}\emptyset t\mbox{}^{\prime}\equiv t\mbox{}t\mbox{}^{\prime}$.
@@ 6.3.3 N-gram Language Models
where, in turn, the probability of a sentence $S\mbox{}_{i}=t\mbox{}_{i,1}t\mbox{}_{i,2}\cdots t\mbox{}_{i,z_{i}}$ is:
$P(S\mbox{}_{i})=\prod_{j=1}^{z_{i}+1}P(t\mbox{}_{i,j}|t\mbox{}_{i,j-N+1}t\mbox{}_{i,j-N+2}\cdots t\mbox{}_{i,j-1})$ (14)
and we make the conventions $t\mbox{}_{i,z_{i}+1}=\langle\text{EOS}\rangle$, $t\mbox{}_{i,-1}=t\mbox{}_{i,-2}=\cdots=t\mbox{}_{i,-N+1}=\langle\text{BOS}\rangle$. As a parenthetical remark, we observe that the probability in Eq. (13) is conditional on the number $m$ of sentences that compose the document $\mathcal{D}_{m}$, as reflected by our notation.
@@ 6.3.4 Kneser-Ney Smoothing
The functions $\alpha(\cdot)$ and $\gamma(\cdot)$ are defined by:
@@
Finally, $D(1),\,D(2),\ldots$ are free parameters satisfying $10<D(k)<$ 1, whose optimal values can be estimated from data. The parameter space is usually made finite by assuming $D(k)=D(r)$ for some $r$ and $k\geq r$, in which case we can rewrite the infinite sum in (18) as:
@@
  • As shown in Eq. (16), the $1$-gram distribution is interpolated with the uniform distribution over the extended dictionary $\mathcal{T}\mbox{}^{*}$ (see next point). This is relevant to our application, since documents featuring tokens not seen during training would otherwise be assigned zero probability, i.e. infinite perplexity.
@@ 6.4 The Likelihood Ratio Framework
  • The Defense Hypothesis ($H_{d}$): The disputed document ($\mathcal{D}_{\mathcal{U}}$) was not written by the author of the known document(s) (the defendant, $\mathcal{A}$), but was written by someone else, $\neg\mathcal{A}$
@@
It is important to note that the forensic expert cannot estimate and legally must not refer to the posterior odds. They cannot estimate it logically because they are not privy of the trier-of-fact’s belief regarding the hypotheses and do not have access to other pieces of evidence presented prior to the linguistic text evidence. Legally, they must not refer to it because the posterior odds represents the ultimate issue of being guilty or innocence [88, p. 96] and this determination is a specific responsibility reserved only for the trier-of-fact. The odds form of Bayes’ Theorem explicitly outlines the role of the forensic scientist. In the Likelihood Ratio framework, their responsibility lies in estimating the strength of evidence in the form of a likelihood ratio. As indicated by the rightmost term of Equation 21, the likelihood ratio is the ratio between the probability of the evidence given $H_{p}$ and the probability of the same evidence given $H_{d}$. The Likelihood Ratio serves as a gradient, quantifying how much more likely the evidence is given one hypothesis over the other [44]. When the LR is greater than one, it provides stronger support for $H_{p}$, whereas when it is smaller than one, it offers stronger support for $H_{d}$. The LR framework is now accepted as a logically and legally correct framework for interpreting forensic evidential analyses [3, 108], having been studied for various evidential types [21, 22, 36, 92, 96, 138]. The LR framework has also been supported by the relevant scientific and professional associations [2, 99, 30, 105, 9, 43].
@@ 6.4.1 The Log-Likelihood Ratio Cost ($C_{\text{llr}}\mbox{}$)
In numerous source-identification tasks, the system’s performance is commonly evaluated in terms of identification Accuracy and/or identification error rate. Metrics such as precision, recall, and equal error rate (EER) are typical in this regard. While these metrics offer a quick overview of performance and are intuitively easy to comprehend, they are not suitable for assessing likelihood ratio-based inference systems [93, p. 93].
@@
Primarily based on categorical decisions with specific thresholds, these metrics are inadequate for evaluating such systems. Firstly, metrics relying on categorical thresholding implicitly assess how effectively the system correctly identifies the source of evidence. In other words, these metrics implicitly pertain to the strength of the hypothesis rather than the strength of the evidence [93, p. 93]. Secondly, categorical metrics fail to capture the gradient nature of likelihood ratios [93, p. 93]. For instance, in a same-author comparison, if one system returns $\log_{10}(\text{\emph{LR}})=3$ while another yields $\log_{10}(\text{\emph{LR}})=2$, the former is superior as it supports the correct hypothesis more strongly. However, using a categorical metric would assess both systems in the same way, disregarding the differing strengths of evidence supporting the correct hypothesis.
@@
The conventional metric employed for evaluating the performance of a system within the Likelihood Ratio Framework is the log-likelihood ratio cost ($C_{\text{llr}}\mbox{}$). The computation of the $C_{\text{llr}}\mbox{}$ involves utilizing a set of LRs for same-source scenarios (e. g., same-authorship, Y) and a set for different-source scenarios (e. g., different-authorship, N). The computation is performed using the following equation [26, 104, 131]
@@
$C_{\text{llr}}\mbox{}=\frac{1}{2}\left(\frac{1}{\gamma}\sum_{i}^{\gamma}\log_{2}\left(1+\frac{1}{\emph{LR}_{i}}\right)+\frac{1}{\eta}\sum_{j}^{\eta}\log_{2}\bigg{(}1+\emph{LR}_{j}\bigg{)}\right)$ (22)
[…]
@@
$C_{\text{llr}}\mbox{}$ serves as an assessment metric for the overall performance of an LR-based inference system and can be decomposed into two distinct metrics: $C_{\text{llr}}^{\text{min}}\mbox{}$ and $C_{\text{llr}}^{\text{cal}}\mbox{}$. The former evaluates the discrimination performance and the latter the calibration performance of the system. The $C_{\text{llr}}^{\text{min}}\mbox{}$ is determined by computing the losses for the optimized LRs through the non-parametric pool-adjacent-violators algorithm. The disparity between $C_{\text{llr}}\mbox{}$ and $C_{\text{llr}}^{\text{min}}\mbox{}$ is expressed as $C_{\text{llr}}^{\text{cal}}\mbox{}$; in other words, $C_{\text{llr}}\mbox{}$ equals the sum of discrimination loss ($C_{\text{llr}}^{\text{min}}\mbox{}$) and calibration loss ($C_{\text{llr}}^{\text{cal}}\mbox{}$). More in-depth descriptions of $C_{\text{llr}}\mbox{}$, encompassing both $C_{\text{llr}}^{\text{min}}\mbox{}$ and $C_{\text{llr}}^{\text{cal}}\mbox{}$, can be found in work by Brümmer and du Preez [26], Drygajlo et al. [41], and Meuwly et al. [90]. The $C_{\text{llr}}\mbox{}$ has already been effectively utilized for evaluating LR-based forensic text comparison systems [62, 63, 64, 97].
6.5 Baseline AV Methods
For the evaluation of LambdaG, we selected seven baseline AV methods that were covered in our literature review (see Section 1). Specifically, we chose IM [84], COAV [57], TAVeer [55], FeVecDiff (LR [133] + MLP [132]) and ADHOMINEM [19]. The selection of these methods was guided by several key considerations. ADHOMINEM, FeVecDiff (LR + MLP) and TAVeer were the three top approaches in the PAN AV 2020 competition [70]. Furthermore, FeVecDiff (LR) ranked third in the PAN AV 2021 competition [71], closely followed by the second-placed method. Regarding IM and COAV, Halvani [51] demonstrated that, across ten different corpora, IM achieved the highest average performance, while COAV ranked second (both in terms of Accuracy and AUC) among twelve evaluated AV methods. COAV has also attracted attention in the AV community. In independent studies, Halvani et al. [57] and Bevendorff et al. [15] came to the conclusion that COAV is on par with Bagnall’s well-known RNN-based AV method [12], the winning approach of the PAN AV 2015 competition [120]. Moreover, COAV was selected as one of six baseline methods in the latest PAN AV 2024 competition, where it outperformed four of them [16, Table 5]. In addition to these six (published) baselines, we have also decided to implement an additional baseline that is used as a common conceptual AV approach in the literature (e. g., [5, 135]). This baseline is described in Section 6.6. Note that with respect to COAV, TAVeer, FeVecDiff (LR + MLP) and ADHOMINEM, we used the original implementations [^11] The original implementations were downloaded from the following Github repositories:
ADHOMINEM $\rightarrow$ https://github.com/boenninghoff/AdHominem
FeVecDiff (MLP) $\rightarrow$ https://github.com/janithnw/pan2020_authorship_verification
FeVecDiff (LR) $\rightarrow$ https://github.com/janithnw/pan2021_authorship_verification
of the authors. Regarding IM, we have used the same implementation as used in [51]. The hyperparameter optimization for FeVecDiff (LR + MLP) and ADHOMINEM was performed by splitting each training corpus $\mathcal{C_{\,\textrm{train}}}$ into a training subset $\mathcal{C_{\,\textrm{train}*}}$ and a validation set $\mathcal{C_{\,\textrm{val}}}$, following an 80/20 ratio.
6.6 SiamBERT
As briefly explained in the previous section, we decided to add another baseline method to our evaluation, which we refer to as SiamBERT. For this purpose, we have implemented an AV method based on a Siamese BERT neural network with a structure similar to the well-known sentence BERT architecture [106]. SiamBERT has a similar functionality to the method of Ishihara et al. [66], in which likelihood-ratios are estimated via cosine similarity and probabilistic linear discrimination analysis.
SiamBERT is based on the BERT-Large, Cased [^12] https://huggingface.co/google-bert/bert-large-cased model with the corresponding tokenizer [37], which are pre-trained on English texts and take into account character-sensitive capitalization. For a given document $\mathcal{D}$, we first tokenize it into a sequence of tokens. Since our BERT model is constrained to processing a maximum of 512 tokens per input, we segment the token sequence into a set of adjacent chunks $\varsigma_{1},\ldots,\varsigma_{m}$, where each chunk $\varsigma_{i}$ consists of 512 tokens, except for the final chunk, which may contain fewer. These chunks are then embedded as feature vectors, by encoding each $\varsigma_{i}$ with the BERT model and taking the average of the resulting token-level embedding vectors over the tokens in a chunk. Finally, we obtain the document’s overall representation $\mathcal{F}\mbox{}$ by averaging these $m$ chunk-level feature vectors, embedding the document $\mathcal{D}$ in a continuous vector space. Given a verification case $c=(\mathcal{D}_{\mathcal{U}},\mathcal{D}_{\mathcal{A}})$, we first embed both documents into their corresponding feature vectors $\mathcal{F}\mbox{}_{\mathcal{U}}$ and $\mathcal{F}\mbox{}_{\mathcal{A}}$ (as described above) and then compute their cosine similarity $s$. Finally, $c$ is classified as either Y (i. e., $\mathcal{A}=\mathcal{U}$) if $s>\theta$, otherwise as N (i. e., $\mathcal{A}\neq\mathcal{U}$), where $\theta$ is a predefined threshold.
To train SiamBERT, we first split the training corpus $\mathcal{C_{\,\textrm{train}}}$ into $\mathcal{C_{\,\textrm{train}*}}$ and $\mathcal{C_{\,\textrm{val}}}$ based on an 80/20 ratio to optimize the hyperparameters. Based on $\mathcal{C_{\,\textrm{train}*}}$, we then train the model parameters using the MSE loss function (calculating the mean squared error between the cosine similarity and the binary label, where 1 represents a Y-case), the AdamW optimizer, a batch size of $4$ and a learning rate of $0.00005$ for $100$ epochs (scheduled with a linear warm up in the first $10$ epochs and a cosine decay to zero in the remaining). To optimize the hyperparameter for the epoch count, we evaluated the AUC at each epoch on $\mathcal{C_{\,\textrm{val}}}$ and selected the model with the highest AUC. Using this selected model, we then computed the cosine similarities for the verification cases in $\mathcal{C_{\,\textrm{train}}}$ and determined the optimal threshold $\theta$ that maximized Accuracy on $\mathcal{C_{\,\textrm{train}}}$.
7 Qualitative Data Exploration
This Section contains a series of examples of how an analyst can visualize the results of LambdaG. To aid readability, we replaced the original placeholders used in the POSNoise paper (see [53, Table 2]) within the visualizations with their corresponding POS tags, according to the following mapping:
[…]